OpenStudy (clongoria2):

Help please! Medal and fan Part A: Divide (9x4y3 + 3x3y2 - 6x2y - 12x2y4) by -3x2y. Show your work, and justify each step. Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? Part C: What is the degree and classification of the polynomial you got in Part A?

OpenStudy (clongoria2):

If you have the time i would also be grateful if you explain the answer.

OpenStudy (anonymous):

I will help u

OpenStudy (anonymous):

IGreen can u help her plz

OpenStudy (igreen):

Divide all the whole numbers by -3

OpenStudy (igreen):

Subtract all the exponents of 'x' by 2, and subtract all the exponents of 'y' by 1.

OpenStudy (igreen):

\(\dfrac{9x^4y^3 + 3x^3y^2 - 6x^2y - 12x^2y^4}{-3x^2y} \) \(\dfrac{9x^4y^3}{-3x^2y} \rightarrow \dfrac{9}{-3} x^{4-2}y^{3-1}\) Simplify that

OpenStudy (igreen):

Divide 9 / -3, subtract 4 - 2, and subtract 3 - 1

OpenStudy (clongoria2):

-3x^2^2 + -1xy + -2xy + -4xy^3

OpenStudy (clongoria2):

This is what i got

OpenStudy (clongoria2):

-3x^2y^2

OpenStudy (igreen):

This is correct: "-3x^2^2 + -1xy"

OpenStudy (igreen):

-2xy is incorrect..

OpenStudy (igreen):

So is -4xy^3

OpenStudy (igreen):

\(\dfrac{-6x^2y}{-3x^2y}\)

OpenStudy (igreen):

The variables cancel out because they're the same..so just divide -6 / -3, what do you get?

OpenStudy (clongoria2):

Oh i see its because the negatives cancel each other out. So -6x^2y / -3x^2y = 2xy?

OpenStudy (clongoria2):

@iGreen